OpenStudy (anonymous):

Henry grabs one shoe at a time at random without replacement from a box of ten shoes in his closet. The box contains 2 brown shoes, 4 black shoes, and 4 sneakers. What is the probability that the first two shoes he chooses will be brown? Round to the nearest percent if necessary.

OpenStudy (anonymous):

1/50

OpenStudy (anonymous):

Is it 20%? The answers are A)2% B)20% C)22% D)40%

OpenStudy (anonymous):

2%

OpenStudy (anonymous):

I don't understand can you please explain how you got to that answer?

OpenStudy (anonymous):

First take, 2/10 = 1/5 = .2 Second take, 1/10 = .1 .2 * .1 = .02 .02 = 2%

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

without replacement!!!! the denominator changes!!!!

OpenStudy (anonymous):

(2/10)*(1/9)

OpenStudy (anonymous):

not .02

OpenStudy (anonymous):

... my answer is still correct, round XD

OpenStudy (anonymous):

.022222222222

OpenStudy (anonymous):

yeah but you r method is dead wrong... so how could he do it again?

OpenStudy (anonymous):

First take, 2/10 = 1/5 = .2 Second take, 1/9 = .11111111111111111111111111111111 .2 * .111111111111111111111111 = .02222222222222222222222 .02222222222222222222222222 = (rounded) 2%

OpenStudy (anonymous):

wow... oblivious.

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

It was correct anyways so whatever